### 29 Apr 2026 v2

# Don’t Retrieve, Navigate: Distilling Enterprise Knowledge into Navigable Agent Skills for QA and RAG

Abstract
Retrieval-Augmented Generation (RAG) grounds LLM responses in external evidence but treats the model as a passive consumer of search results: it never sees how the corpus is organized or what it has not yet retrieved, limiting its ability to backtrack or combine scattered evidence. We present Corpus2Skill, which distills a document corpus into a hierarchical skill directory offline and lets an LLM agent navigate it at serve time. The compilation pipeline iteratively clusters documents, generates LLM-written summaries at each level, and materializes the result as a tree of navigable skill files. At serve time, the agent receives a bird’s-eye view of the corpus, drills into topic branches via progressively finer summaries, and retrieves full documents by ID. Because the hierarchy is explicitly visible, the agent can reason about where to look, backtrack from unproductive paths, and combine evidence across branches. On WixQA, an enterprise customer-support benchmark for RAG, Corpus2Skill outperforms dense retrieval, RAPTOR, and agentic RAG baselines across all quality metrics. We further evaluate generalization on nine RAGBench subsets reformulated as retrieval-stress benchmarks: Corpus2Skill attains the highest macro-average F1 across the full 10-dataset suite and characterizes a clear regime — single-domain, atomic-document corpora — where corpus navigation is the right primitive, while flat retrieval remains preferable for open-domain or extractive pools.
Retrieval-Augmented Generation, Agent Skills, Hierarchical Clustering, Knowledge Navigation
Yiqun Sun   Pengfei Wei   Lawrence B. Hsieh
Magellan Technology Research Institute (MTRI)
{duke.sun, pengfei.wei, lawrence.hsieh}@mtri.co.jp
Code: https://github.com/dukesun99/Corpus2Skill
[…]
1 Introduction
Figure 1: Retrieve vs. Navigate. Traditional RAG passively feeds fixed passages to the LLM. Corpus2Skill distills the corpus into a navigable skill hierarchy that the agent actively explores, backtracks, and drills into to locate evidence.
Enterprise knowledge bases contain thousands of heterogeneous documents: support articles, product guides, policy documents, and FAQs (Gao et al., 2024). Large language models alone cannot answer questions about these corpora reliably: they lack access to private, frequently updated content and are prone to hallucinating plausible but incorrect details (Lewis et al., 2020; Gao et al., 2024). Retrieval-Augmented Generation (RAG; Lewis et al., 2020) was introduced to solve these problems by grounding LLM responses in retrieved evidence. By embedding the query, fetching similar passages from a vector index, and conditioning the model’s answer on the retrieved context, RAG enables LLMs to leverage enterprise-specific knowledge they were never trained on, providing factual, up-to-date answers while reducing hallucination. This pipeline has become the dominant paradigm for enterprise QA, customer support, and compliance workflows (Gao et al., 2024; Singh et al., 2025). Yet it presents a structural limitation for more complex queries: the LLM never sees the forest for the trees.
Consider a support agent faced with the query “How do I switch my business from sole proprietorship to LLC on Wix?” This query spans multiple topics: account types, payment provider configuration, and legal entity management, each covered in different sections of the knowledge base. A flat retrieval system returns the top-$k$ passages closest to the query embedding, which are likely to match on surface-level terms (“sole proprietorship,” “LLC”) but miss the operationally critical article explaining that account types cannot be changed directly and require contacting support. The LLM receives these fragments but has no understanding of how the corpus is organized, what topic areas exist beyond the retrieved set, or whether the best-matching documents were even in the retrieval window. It cannot reason about where else to look, whether its evidence is complete, or how the retrieved passages relate to the broader knowledge structure. In short, conventional RAG treats the LLM as a passive consumer of search results, leaving the model’s planning and reasoning capabilities largely unused.
Agentic RAG systems (Yao et al., 2023; Schick et al., 2023) partially address this by allowing the LLM to issue multiple search queries iteratively. But even with tool access, the agent operates without a map: it must guess productive search terms, has no way to survey the corpus structure, and cannot systematically narrow its focus from broad topics to specific documents. Each search query is a shot in the dark, guided only by the agent’s prior knowledge and whatever fragments previous queries returned. For the business-type query above, an agentic retrieval system would need to independently discover that the answer involves both the payments ecosystem and account verification policies — topics whose connection is not apparent from keyword similarity alone.
Hierarchical approaches such as RAPTOR (Sarthi et al., 2024) and GraphRAG (Edge et al., 2024) organize documents into trees or graphs with LLM-generated summaries at internal nodes. These summaries capture thematic structure that flat embeddings miss. However, these systems still rely on embedding-based similarity search at query time: RAPTOR indexes all tree nodes in a flat vector store, while GraphRAG requires a graph database and entity extraction pipeline. The hierarchy enriches the retrieval candidates but remains invisible to the LLM at query time; the model still receives search results without understanding the organizational structure that produced them. More recent work pushes further: StructRAG (Li et al., 2025) restructures retrieved information into tables or graphs at inference time, and BookRAG (Wang et al., 2025) builds hierarchical indices with agent-based planners. Yet these systems still invoke embedding-based search at query time and do not give the agent direct access to the organizational structure itself.
Meanwhile, a parallel line of work on LLM agent skills (Wang et al., 2023; Jiang et al., 2026) has shown that agents perform complex tasks more effectively when given structured, reusable knowledge modules that they can selectively load and follow. Agent platforms now support filesystem-based skill packages [^1] https://docs.anthropic.com/en/docs/agents-and-tools/agent-skills that the agent discovers via metadata and reads on demand, a mechanism known as progressive disclosure (Section 2.3). This suggests an alternative to retrieval: rather than searching for passages, distill the corpus into a navigable skill hierarchy and let the agent explore it.
We propose exactly this: make the hierarchy itself the primary interface for the LLM. Rather than searching over a hierarchy, the agent navigates through it. This is the core idea behind Corpus2Skill (Figure 1): a one-time offline compilation phase transforms a document corpus into a navigable skill tree, and at serve time, the LLM agent browses this tree to find information, much like a human navigating a well-organized file system or table of contents.
We evaluate Corpus2Skill on WixQA (Cohen et al., 2025), an enterprise customer-support benchmark designed for evaluating RAG systems over real-world knowledge bases. Corpus2Skill achieves the highest quality across all metrics, outperforming flat retrieval, RAPTOR, and a multi-turn agentic RAG system, while requiring no embedding index or vector database at serve time. We complement this anchor result with a generalization study on nine RAGBench (Friel et al., 2024) subsets reformulated as retrieval benchmarks: Corpus2Skill attains the highest macro-average F1 over the full ten-dataset benchmark, and the per-dataset dispersion isolates the regime where navigation is the right primitive (single-domain, atomic-document corpora) versus the regime where flat retrieval remains preferable (open-domain abstracts, very long extractive documents). These results suggest a complementary design point for knowledge-grounded systems: invest computation at compile time to make the corpus navigable, rather than solely at query time to make it searchable.
We summarize our contributions as follows.
  1. 1.
    We identify corpus navigability as an underexplored dimension in current RAG systems: LLMs lack a structured overview of the knowledge base, which limits their ability to systematically locate the best evidence.
  2. 2.
    We propose Corpus2Skill, a compile-then-navigate framework: an offline pipeline transforms a raw corpus into a hierarchical skill directory via iterative clustering and LLM-based summarization, and at serve time an LLM agent navigates this hierarchy through file browsing and document lookup tools, requiring no embedding index or vector database at serve time.
  3. 3.
    We evaluate on WixQA (Cohen et al., 2025) against five baselines, conduct ablation studies on hierarchy structure, exploration budget, and serving model choice, and run a 9-subset generalization study on RAGBench (Friel et al., 2024) that characterizes which corpus shapes Corpus2Skill helps and which it does not.
2 Background: Skills in LLM Agents
2.1 From Tools to Skills
Modern LLM agents interact with external systems through tools: atomic, stateless functions such as web search or database queries (Schick et al., 2023; Yao et al., 2023). Tools give an agent the ability to act but do not tell it how to approach a complex task. Agent skills fill this gap: reusable modules that package procedural knowledge or multi-step workflows into a form the agent can load and follow (Jiang et al., 2026). Unlike a tool, a skill shapes reasoning over multiple steps, persisting across tool calls to guide the agent’s decisions throughout a task. This distinction is increasingly important as agents tackle longer-horizon tasks where choosing the right sequence of tool calls requires domain-specific strategy, not just tool availability.
Formally, Jiang et al. (2026) define a skill as a tuple $S=(C,\pi,T,R)$: an applicability condition $C$, an execution policy $\pi$, a termination condition $T$, and a reusable interface $R$. This distinguishes skills from one-off plans (no reusable interface), episodic memories (no execution policy), and atomic tools (no multi-step structure).
2.2 Skill Representations
Skills take several forms in practice:
  • Code-as-skill. Voyager (Wang et al., 2023) stores successful action sequences as executable functions in a growing library.
  • Natural-language playbooks. SkillX (Wang et al., 2026) and EvoSkill (Alzubi et al., 2026) distill agent experience into structured textual instructions.
  • Filesystem-based packages. A directory containing a structured instruction file (SKILL.md) alongside optional scripts and data assets. This is the pattern adopted by recent agent platforms and the one Corpus2Skill builds on.
Code-based skills offer precision but require execution environments; natural-language skills are portable and interpretable at the cost of ambiguity. Filesystem-based packages balance the two by combining structured text with auxiliary resources that the agent can access on demand.
2.3 Progressive Disclosure
Loading all skill content at startup would overwhelm the context window; loading nothing would leave the agent unable to choose. The filesystem pattern resolves this via progressive disclosure (Shneiderman, 1996):
  1. 1.
    Metadata (at startup): skill name and one-line description (<<100 tokens each).
  2. 2.
    Instructions (on demand): the full SKILL.md is read when the agent selects a skill.
  3. 3.
    Resources (as needed): supporting files remain on disk, consuming zero tokens until accessed.
The agent thus maintains awareness of many skills while paying the token cost only for those actually used.
2.4 From Procedural Skills to Informational Skills
In prior work, skills encode how to perform a task: workflows, decision trees, or code patterns. Corpus2Skill repurposes the same mechanism to encode what a corpus contains. Each skill directory represents a topical partition, with SKILL.md and INDEX.md files that summarize its contents at increasing levels of detail. The agent follows the same progressive disclosure workflow (description $\to$ SKILL.md $\to$ INDEX.md $\to$ document) but navigates an information hierarchy rather than executing procedural instructions. This reinterpretation—from procedural to informational skills—is the central design insight of Corpus2Skill.
3 Corpus2Skill Framework
Corpus2Skill operates in two phases: a one-time offline compile phase that transforms a raw document corpus into a browsable skill tree, and a per-query online serve phase where an LLM agent navigates this tree with full visibility into the corpus structure. Figure 2 illustrates the full architecture.
3.1 Problem Setup
Let $\mathcal{D}=\{d_{1},\dots,d_{N}\}$ be a corpus of $N$ documents. Given a natural-language query $q$, the task is to produce an answer $a$ grounded in a subset $\mathcal{D}_{q}\subseteq\mathcal{D}$ of relevant documents. Standard RAG systems select $\mathcal{D}_{q}$ via a retrieval function $f(q,\mathcal{D})\to\mathcal{D}_{q}$ that operates as a black box from the LLM’s perspective: the model receives $\mathcal{D}_{q}$ but has no visibility into what $\mathcal{D}$ contains or why $\mathcal{D}_{q}$ was chosen.
Corpus2Skill replaces this opaque retrieval step with agent-driven navigation. The compile phase constructs a hierarchical skill forest $\mathcal{S}=\{s_{1},\dots,s_{K}\}$ where each skill $s_{k}$ is a navigable directory tree of summaries covering a topical partition of $\mathcal{D}$. Following the agentic skills formalism of Jiang et al. (2026), each skill can be viewed as a tuple $(C,\pi,T,R)$: the applicability condition $C$ is the skill’s summary (the agent reads it to decide relevance); the execution policy $\pi$ is the navigation workflow (descend through summaries, then retrieve documents); the termination condition $T$ is the agent’s judgment that sufficient evidence has been gathered; and the reusable interface $R$ is the standardized file structure (SKILL.md, INDEX.md) and the get_document tool. At serve time, the agent selects and navigates skills autonomously, making the document selection process transparent and steerable.
Figure 2: Corpus2Skill system architecture. The compile phase (top) embeds documents, builds a multi-level cluster hierarchy with LLM summarization at each level, and materializes the result as a forest of navigable skills. Root clusters map to skill directories (SKILL.md), sub-clusters map to group directories (INDEX.md), and leaf documents are stored in a separate document store. The serve phase (bottom) uses an LLM agent with pre-compiled skills and two tools: file navigation (code_execution) and document lookup (get_document).
3.2 Compile Phase
The compilation pipeline consists of four stages (implementation details and full prompts in Appendices A and C).
Document loading and embedding.
The compiler accepts directories containing standard document formats (.md, .txt, .json, .jsonl). Each document is assigned a content-hash-based unique identifier, ensuring deterministic, collision-free IDs that are independent of the source system’s naming scheme. Documents are truncated to a maximum character length and embedded using a sentence embedding model (Reimers and Gurevych, 2019), producing a dense vector for each document. These embeddings capture semantic similarity between documents and serve as the foundation for the subsequent clustering step.
Iterative hierarchical clustering.
We build a multi-level hierarchy through an iterative bottom-up process controlled by two parameters: the branching ratio $p$ (target number of children per cluster) and the maximum number of top-level clusters $K$. At each level, the current set of $n$ embeddings is partitioned into $\lceil n/p\rceil$ groups via K-Means (Lloyd, 1982) on L2-normalized vectors. Each resulting cluster is then summarized by an LLM that reads representative member texts and produces a concise description of the cluster’s topical coverage, the types of questions the documents answer, and key terms or features mentioned. These summaries are re-embedded using the same sentence embedding model, and the resulting embeddings become the input to the next round of K-Means. This process repeats until the number of remaining clusters falls below $K$, yielding a set of root clusters that form the top level of the hierarchy.
This iterative embed-cluster-summarize loop is inspired by the recursive tree construction in RAPTOR (Sarthi et al., 2024), but adapted for our navigability goal. We use K-Means with hard assignments so that each document maps to exactly one path in the directory tree, and we materialize the hierarchy as browsable files rather than indexing it in a flat vector store. These two choices, deterministic partitioning and filesystem materialization, are what enable agent-driven navigation at serve time. Small clusters below a minimum size are merged into the nearest centroid to avoid orphan groups.
Labeling.
After the clustering hierarchy stabilizes, the LLM generates a short (2 to 5 word) filesystem-safe label for each non-leaf node. These labels serve a dual purpose: they provide human-readable directory names for inspection (e.g., wix-commerce-monetization) and, more importantly, they give the agent semantically meaningful entry points for topic routing. When the agent reads a SKILL.md listing subgroups such as “wix-payments-ecosystem” or “domain-dns-management,” the labels alone often suffice to identify the relevant branch without reading full summaries.
Skill tree construction.
The hierarchy is materialized into a filesystem structure designed for agent navigation:
  • Root clusters become top-level skill directories, each containing a SKILL.md with the cluster summary and a listing of child groups.
  • Sub-clusters become subdirectories with INDEX.md files that list either further subgroups or leaf document identifiers alongside brief summaries of each document.
  • Full document text is stored externally in a documents.json file, referenced by document ID from within the navigation files.
This separation between navigation metadata and document content is a deliberate architectural choice. Navigation files (SKILL.md, INDEX.md) remain small (typically under 2 KB each), containing only summaries and document identifiers. This keeps the agent’s per-step token cost low: reading an INDEX.md to survey 20 documents costs far fewer tokens than reading the documents themselves. Full document content is accessed only when the agent has narrowed its search, via a dedicated get_document tool call. The result is a forest of $K$ independent skill trees, each representing a topical domain within the corpus. The forest structure allows the agent to reason about topic relevance at the coarsest level before committing to a branch. [^2] Anthropic’s current Skills API imposes limits of 8 skills per request, 200 files and 30 MB per skill. These constraints informed our default parameters but are not fundamental to the framework design.
Skill file design.
Each SKILL.md and INDEX.md follows a uniform template with YAML frontmatter (name, description, level, num_documents) that the Skills API reads for progressive disclosure. Compilation prompts produce routing-oriented summaries: the LLM identifies the topic area, question types, and key terms, so that each file functions as a decision point for the navigating agent. Each file also embeds actionable next-step instructions (e.g., “Use get_document with the doc_id”), so the agent always knows which tool to invoke next.
Figure 3 shows abridged examples of both file types. The SKILL.md (top) gives the agent a single-glance overview of all subgroups within a skill; the leaf INDEX.md (bottom) lists individual document IDs with concise titles for targeted retrieval. Unabridged versions at all three hierarchy levels are in Appendix B.
SKILL.md (root level, abridged) ---
name: skill-01-wix-commerce-monetization
description: >
Full Wix commerce and monetization
ecosystem, from storefronts to payments.
level: 3
num_documents: 1513
## Contents
### Sub-groups (directories)
- group-00-wix-restaurants/ (159 docs):
Wix Restaurants orders, menus...
- group-01-wix-payments-ecosystem/ (87):
Payment processing, account mgmt...
- group-02-wix-pos-systems/ (110 docs)
[...16 subgroups total]
INDEX.md (leaf level, verbatim) ---
name: L1-C298
description: >
Fraud prevention and payment security
within Wix Payments.
level: 1 num_documents: 5
### Documents (5 items)
- 5̀6ae76:̀ Best Practices for Accepting
Card Payments
- 4̀48d79:̀ About the ‘Liability Shift’
- b̀b73d9:̀ Declined Payment Due to
Suspected Fraud
- c̀3b01a:̀ About Suspicious Payments
- d̀22b8c:̀ Double Charge on Statement
Figure 3: Abridged SKILL.md (top) and leaf INDEX.md (bottom) from the WixQA compilation. The SKILL.md provides a routing overview of 1,513 documents; the INDEX.md lists individual document IDs with titles for targeted retrieval.
Navigation complexity.
The hierarchy depth grows as $O(\log_{p}N)$: with branching ratio $p$ and $N$ documents, the agent traverses at most $L=\lceil\log_{p}N\rceil$ levels to reach a leaf. At each level $\ell$, the agent reads one navigation file summarizing $O(p)$ children. A single decision at level $\ell$ reduces the candidate set from $N/p^{\ell-1}$ to $N/p^{\ell}$, yielding a multiplicative information funnel:
$N\;\xrightarrow{\div\,p}\;\frac{N}{p}\;\xrightarrow{\div\,p}\;\frac{N}{p^{2}}\;\cdots\;\xrightarrow{\div\,p}\;\frac{N}{p^{L}}\approx 1$ (1)
The total summaries inspected scale as $O(p\cdot\log_{p}N)$ rather than $O(N)$. For WixQA (N=6,221N{=}6{,}221, p=10p{=}10), this gives L=3L{=}3 levels and approximately 30 summaries to locate a single document among thousands. Note that this bound counts summaries inspected under an idealized single-path traversal; in practice, the actual token cost is higher because the Skills API includes previously loaded navigation files in each subsequent call, and the agent may explore multiple branches. The per-query costs reported in Section 4 reflect these real-world overheads. A 100,000-document corpus with p=10p{=}10 would require only one additional level (L=5L{=}5).
3.3 Serve Phase
At serve time, the pre-compiled skill directories are uploaded to Anthropic’s Skills API. The API implements progressive disclosure: skill names and one-line descriptions are pre-loaded into the agent’s context, but full file content is loaded only when the agent explicitly reads a file. This design is critical for token efficiency. Pre-loading all navigation files would flood the context with thousands of tokens of irrelevant summaries; loading nothing would leave the agent blind to the corpus. Progressive disclosure strikes a balance: the agent sees just enough (skill names and descriptions, approximately 200 tokens total) to make an informed first decision, then incrementally loads detail as it descends.
Tools.
The agent receives two tools: (1) code_execution, which allows browsing SKILL.md and INDEX.md files through view commands, providing the agent with full visibility into the hierarchy; and (2) get_document(doc_id), a custom tool that retrieves the full text of a document from the document store given its identifier. The two-tool design reflects the two types of information the agent needs: navigational (where to look) and evidential (what to cite). Keeping them separate means the agent can survey many document summaries cheaply before committing to the token cost of reading full documents.
Navigation workflow.
Because skill descriptions are pre-loaded, the agent begins each query with a bird’s-eye view of the corpus, knowing what topics exist and roughly how many documents each topic covers. A typical query proceeds in 2 to 3 turns:
  1. 1.
    The agent identifies the most relevant skill from pre-loaded descriptions and reads its SKILL.md to understand the subgroup structure.
  2. 2.
    The agent navigates into the relevant subgroup’s INDEX.md, which lists document identifiers with brief titles and summaries.
  3. 3.
    The agent calls get_document to retrieve the full text of one or more promising documents and synthesizes a grounded answer.
At each step, the agent makes an informed decision guided by summaries describing each branch. If a SKILL.md lists ten subgroups and the agent has only explored one, it knows nine remain—an awareness that flat retrieval cannot provide. This structural visibility enables two behaviors: targeted backtracking (abandoning an unpromising branch for a more relevant one) and cross-branch synthesis (combining evidence from multiple subgroups within or across skills). The system prompt enforces a grounding rule (every claim must trace to a retrieved document, not a summary file) and a concise-answer format (word limits, no preamble); see Appendix D.
4 Experiments
4.1 Setup
Dataset.
We evaluate on WixQA (Cohen et al., 2025), a customer support QA benchmark built from the Wix knowledge base. The corpus contains 6,221 support articles covering website building, e-commerce, SEO, marketing, and platform features. The evaluation set consists of 200 expert-written questions with gold-standard answers and gold article IDs.
Compilation.
We compile the WixQA corpus with branching ratio p=10p{=}10 and maximum top-level clusters K=7K{=}7. This produces a 3-level hierarchy with 6 top-level skills, 665 navigation files, and a 13 MB document store. Compilation takes 6.5 minutes on a 32-CPU server.
Baselines.
All methods use the same LLM for answer generation and the same evaluation protocol. We compare five baselines spanning three retrieval paradigms: BM25 (sparse keyword retrieval (Robertson and Zaragoza, 2009)), Dense (Qwen3-Embedding-0.6B (Zhang et al., 2025) with a FAISS index (Johnson et al., 2019)), Hybrid (Reciprocal Rank Fusion (Cormack et al., 2009) of BM25 and Dense), RAPTOR (hierarchical tree retrieval (Sarthi et al., 2024) with UMAP+GMM+BIC clustering), and Agentic (an LLM agent with iterative access to BM25, Dense, and Hybrid tools). The first three are single-shot (top-5 passages, one LLM call); RAPTOR uses collapsed tree retrieval; Agentic is allowed up to 10 interaction rounds. Full implementation details are in Appendix H.
Metrics.
We report four lexical overlap metrics: Token F1 (harmonic mean of token-level precision and recall against the gold answer), BLEU (Papineni et al., 2002) (n-gram precision with brevity penalty), ROUGE-1 and ROUGE-2 (Lin, 2004) (unigram and bigram recall). We additionally report two LLM-judged metrics following the LLM-as-judge paradigm (Zheng et al., 2023), inspired by RAGAS (Es et al., 2024): Factuality measures how factually correct the generated answer is relative to the gold answer (1–5 scale, normalized to 0–1), and Context Recall measures how well the retrieved context covers the essential claims in the gold answer (same scale). We also report input token usage and per-query cost. Details on metric computation and fairness standardization across methods are in Appendix I.
4.2 Main Results
[…]
Method Quality Metrics Cost
[…]
BM25 0.342 0.060 0.364 0.119 0.470 0.386 1.00 708 0.007
Dense 0.363 0.074 0.382 0.141 0.536 0.450 1.00 656 0.008
Hybrid 0.360 0.071 0.380 0.137 0.524 0.410 1.00 698 0.008
RAPTOR 0.389 0.109 0.406 0.189 0.675 0.616 1.00 1,807 0.012
Agentic 0.381 0.098 0.395 0.183 0.719 0.498 5.21 15,483 0.088
Corpus2Skill 0.468 0.143 0.483 0.238 0.739 0.673 2.43 12,601 0.089
Quality.
Table 1 presents the main results. Corpus2Skill achieves the highest scores across all six quality metrics. On Token F1, Corpus2Skill scores 0.468, a 23% relative improvement over Agentic (0.381) and 29% over Dense retrieval (0.363). On Factuality, Corpus2Skill (0.739) exceeds both Agentic (0.719) and RAPTOR (0.675). On Context Recall, Corpus2Skill leads by a clear margin (0.673 vs. 0.616 for RAPTOR and 0.498 for Agentic), indicating that the navigation-based approach surfaces more relevant content than either vector search or multi-turn agentic retrieval.
Hierarchical methods outperform flat retrieval.
Both RAPTOR and Corpus2Skill substantially outperform the three flat retrieval baselines (BM25, Dense, Hybrid), confirming that hierarchical document organization improves answer quality. RAPTOR’s collapsed tree retrieval achieves Factuality of 0.675 and Context Recall of 0.616, substantially above flat methods (Dense: 0.536/0.450, BM25: 0.470/0.386), and Corpus2Skill exceeds RAPTOR by a further 10% on Factuality (0.739) and 9% on Context Recall (0.673). Cluster summaries surface thematic connections that flat embeddings miss, allowing the retrieval system to match queries to relevant document groups rather than individual passages.
Cost analysis.
[…]
4.3 Case Study: Navigation in Action
We present two representative traces showing how Corpus2Skill navigates the hierarchy (Figures 4 and 5).
Query: “I need to switch my business type from sole prop to LLC in order to use an EIN.” Step 1: view wix-commerce-monetization/SKILL.md -> 10 subgroups; picks "wix-payments- ecosystem" (87 docs) Step 2: view wix-payments-ecosystem/INDEX.md -> 8 leaf groups; picks group covering account management Step 3: view group-05/INDEX.md -> doc d56cc79: "Changing Your Wix Payments Account Type" Step 4: get_document("d56cc797742e") -> Retrieves full article Answer: "You cannot change your account type directly. Contact Wix Customer Care."
Figure 4: Trace 1: Direct navigation. The agent descends through a single skill branch to find the relevant document in 4 steps.
Query: “I want to change the currency for my course.” Step 1: view wix-commerce-monetization/SKILL.md -> 15 subgroups; "wix-online-programs" and "wix-billing-documents" relevant Step 2: view wix-online-programs/INDEX.md -> 5 leaf groups; picks group covering program settings (18 docs) Step 3: view group-03/INDEX.md -> doc 7a9dc6: "Online Programs: Editing Your Program Settings" Step 4: get_document("7a9dc6f2de90") -> Program settings doc; mentions currency is a site-level setting Step 5: view wix-billing-documents/INDEX.md -> picks group covering currency management (18 docs) Step 6: view group-00/INDEX.md -> doc 29139b: "Setting Your Currency for Accepting Payments" Step 7: get_document("29139bd7c379") Answer: "Go to Settings > Language & Region > Currency. This site-level setting applies to all Wix products including Online Programs."
Figure 5: Trace 2: Cross-branch navigation. The agent visits two subgroups within the same skill to combine evidence from different topic branches.
Patterns.
These traces illustrate two common navigation patterns: direct descent (Figure 4: skill $\to$ subgroup $\to$ leaf $\to$ document) and cross-branch exploration (Figure 5: the agent visits multiple subgroups within the same skill to combine evidence). In both cases, the agent’s decisions are guided by summaries at each level. Full traces with raw tool calls are provided in Appendix E.
4.4 Generalization to RAGBench
Table 2: Generalization study on 9 RAGBench subsets, 200 queries each (or all available when smaller), evaluated on the expanded retrieval pool (union of train+val+test documents, deduplicated by content hash). All methods use Claude Sonnet 4.6 with Anthropic ephemeral prompt caching. Quality cells shaded green (darker = better) and \$/q cells shaded red (darker = higher); the per-dataset best value of each quality metric is shown in bold. Datasets are grouped by the structural pattern that determines whether navigation helps; see Table 3 for the full 10-dataset macro-average. “Inp.” is per-query input tokens billed at the full input rate (in thousands); “Turns” counts outer LLM round-trip API calls only (server-side tool invocations within a single request, e.g. Skills / code-execution skill-file reads, are not counted).
Dataset Method Quality Metrics Cost
[…]
(a) Atomic, single-domain corpora — Corpus2Skill wins
Emanual BM25 0.413 0.108 0.435 0.195 0.658 0.552 1.00 0.54 0.006
Dense 0.423 0.116 0.444 0.205 0.705 0.656 1.00 0.50 0.006
Hybrid 0.429 0.118 0.451 0.209 0.727 0.659 1.00 0.57 0.006
Agentic 0.456 0.147 0.475 0.245 0.836 0.624 3.32 9.0 0.044
Corpus2Skill 0.497 0.160 0.514 0.279 0.794 0.762 3.11 7.2 0.054
DelucionQA BM25 0.413 0.149 0.431 0.231 0.727 0.688 1.00 0.68 0.006
Dense 0.438 0.167 0.458 0.257 0.812 0.740 1.00 0.65 0.006
Hybrid 0.432 0.161 0.451 0.250 0.774 0.708 1.00 0.67 0.006
Agentic 0.441 0.178 0.459 0.272 0.838 0.723 2.83 6.4 0.037
Corpus2Skill 0.462 0.174 0.481 0.285 0.782 0.714 2.71 8.8 0.062
HAGRID BM25 0.325 0.109 0.334 0.195 0.831 0.831 1.00 0.66 0.005
Dense 0.317 0.104 0.326 0.190 0.839 0.881 1.00 0.66 0.005
Hybrid 0.328 0.112 0.337 0.198 0.834 0.858 1.00 0.67 0.005
Agentic 0.365 0.133 0.374 0.223 0.841 0.842 2.31 4.2 0.022
Corpus2Skill 0.417 0.165 0.427 0.262 0.824 0.759 2.17 8.5 0.071
(b) Heterogeneous or table-heavy corpora — flat retrieval competitive, Corpus2Skill close
CovidQA BM25 0.203 0.048 0.212 0.100 0.547 0.548 1.00 0.67 0.006
Dense 0.214 0.049 0.224 0.107 0.643 0.645 1.00 0.64 0.006
Hybrid 0.218 0.053 0.227 0.111 0.623 0.586 1.00 0.66 0.006
Agentic 0.200 0.052 0.206 0.107 0.635 0.514 4.05 10 0.056
Corpus2Skill 0.206 0.044 0.217 0.088 0.452 0.396 9.39 17 0.174
ExpertQA BM25 0.310 0.044 0.337 0.086 0.490 0.433 1.00 0.70 0.007
Dense 0.330 0.050 0.358 0.095 0.580 0.526 1.00 0.71 0.007
Hybrid 0.324 0.049 0.352 0.093 0.542 0.487 1.00 0.71 0.007
Agentic 0.289 0.042 0.310 0.092 0.685 0.435 4.44 14 0.085
Corpus2Skill 0.288 0.039 0.311 0.077 0.448 0.295 2.83 9.0 0.088
TechQA BM25 0.381 0.082 0.406 0.150 0.527 0.310 1.00 0.95 0.009
Dense 0.380 0.087 0.404 0.154 0.565 0.303 1.00 0.97 0.009
Hybrid 0.383 0.083 0.409 0.152 0.567 0.304 1.00 0.96 0.009
Agentic 0.297 0.069 0.313 0.132 0.606 0.291 6.36 28 0.151
Corpus2Skill 0.328 0.067 0.350 0.123 0.424 0.281 4.34 27 0.293
FinQA BM25 0.354 0.108 0.364 0.177 0.341 0.362 1.00 0.73 0.005
Dense 0.401 0.122 0.414 0.206 0.525 0.517 1.00 0.74 0.006
Hybrid 0.390 0.118 0.401 0.198 0.446 0.434 1.00 0.73 0.006
Agentic 0.367 0.114 0.375 0.199 0.566 0.493 4.44 15 0.078
Corpus2Skill 0.382 0.132 0.389 0.217 0.555 0.495 6.25 50 0.491
(c) Homogeneous-tabular or long-extractive corpora — Corpus2Skill loses
TatQA BM25 0.307 0.098 0.317 0.166 0.464 0.468 1.00 0.55 0.005
Dense 0.304 0.086 0.313 0.154 0.519 0.512 1.00 0.63 0.006
Hybrid 0.315 0.094 0.325 0.165 0.537 0.518 1.00 0.59 0.005
Agentic 0.351 0.127 0.359 0.213 0.614 0.546 3.04 8.7 0.045
Corpus2Skill 0.246 0.076 0.253 0.135 0.402 0.363 8.43 80 0.969
CUAD BM25 0.314 0.034 0.338 0.082 0.280 0.191 1.00 0.72 0.007
Dense 0.302 0.033 0.325 0.083 0.282 0.170 1.00 0.70 0.008
Hybrid 0.308 0.034 0.331 0.085 0.277 0.204 1.00 0.71 0.007
Agentic 0.249 0.035 0.265 0.084 0.260 0.193 4.29 13 0.079
Corpus2Skill 0.211 0.028 0.227 0.060 0.229 0.201 8.82 55 0.369
Table 3: Macro-average over the full 10-dataset benchmark (WixQA anchor + 9 RAGBench subsets, Table 2). Per-metric best is in bold. Corpus2Skill leads on every lexical-overlap metric; Agentic wins Factuality (it falls back to “I don’t know” more often, which the judge rewards), and Dense wins Context Recall (which rewards span overlap that embedding retrieval directly optimizes for).
[…]
BM25 0.336 0.084 0.354 0.150 0.534 0.477 1.00 0.006
Dense 0.347 0.089 0.365 0.159 0.601 0.540 1.00 0.007
Hybrid 0.349 0.089 0.367 0.160 0.585 0.517 1.00 0.007
Agentic 0.339 0.099 0.353 0.175 0.660 0.516 4.03 0.068
Corpus2Skill 0.350 0.103 0.378 0.177 0.565 0.494 5.04 0.266
To probe how Corpus2Skill behaves outside its target domain, we run the same five methods on nine subsets of RAGBench (Friel et al., 2024): CovidQA (Möller et al., 2020), DelucionQA (Sadat et al., 2023), Emanual (Nandy et al., 2021), ExpertQA (Malaviya et al., 2024), TechQA (Castelli et al., 2020), FinQA (Chen et al., 2021), CUAD (Hendrycks et al., 2021), HAGRID (Kamalloo et al., 2023), and TatQA (Zhu et al., 2021), sampling 200 test queries per subset (or all available when smaller). RAGBench is published as a generator-evaluation benchmark with per-row gold contexts; to convert it into a retrieval benchmark we build the expanded retrieval pool as the union of all train+val+test documents, deduplicated by content hash. This is intentionally harder than the original RAGBench protocol: every test query must compete against thousands of distractors, and there is no curated retrieval corpus to fall back on. Pool sizes range from 221 (Emanual) to 12,727 (TatQA).
Headline result.
Corpus2Skill attains the highest macro-average F1 (0.350) across the 10-dataset benchmark (Table 3), narrowly ahead of Hybrid (0.349) and Dense (0.347), and it leads on every lexical-overlap metric (BLEU, ROUGE-1, ROUGE-2). Agentic continues to win Factuality (0.660), and Dense wins Context Recall (0.540) because the LLM-judged CtxR rewards span overlap that embedding retrieval is explicitly optimized for. The picture is therefore not a uniform sweep — Corpus2Skill wins on 4 of 10 datasets by a margin of 0.02–0.10 F1, ties on 2, and loses on 4 (Table 2) — but the per-dataset dispersion lets us characterize which corpora navigation actually helps.
When does corpus navigation help?
Three structural properties consistently separate wins from losses:
(i) Atomic, topically-coherent documents. On WixQA, Emanual, DelucionQA, and HAGRID — corpora where each document covers a single feature, entity, or how-to — Corpus2Skill beats every retrieval baseline by 0.02–0.09 F1. HAGRID is particularly informative: even though its corpus is open-domain Wikipedia, every article is by construction a coherent unit about one entity, and the cluster summarizer recovers a clean per-topic taxonomy. The hierarchy gives the agent useful disambiguation that flat retrievers cannot exploit.
(ii) Document granularity in the 0.5–3k character range. The wins cluster around documents long enough to contain a complete answer ($\geq 500$ characters) but short enough that an LLM-written summary captures the article ($\leq 3000$ characters). On CUAD (legal contracts averaging $\approx$7,500 characters of mixed clauses) and TechQA (IBM technotes averaging 3,740 characters), single documents straddle multiple clause types or symptom categories, so any single skill placement is provably suboptimal and Corpus2Skill loses to flat retrieval.
(iii) A recoverable implicit taxonomy. On TatQA, which is essentially homogeneous SEC-style financial-statement tables, the level-1 cluster summaries collapse onto near-identical labels (“corporate financial reporting”) and the agent can no longer distinguish branches at the top level. Corpus2Skill hits max_turns on a third of queries and the cost balloons to \$0.97/query, the worst case in our evaluation. ExpertQA exhibits a milder version of the same failure mode: queries span dozens of unrelated scientific fields, summaries become generic, and the navigator cannot exploit them. Agentic search wins both subsets, presumably because the lack of a useful hierarchy makes the search-then-read pattern strictly more efficient.
Where flat retrieval is the right default.
On CovidQA, the corpus is short, near-duplicate research abstracts whose information density is already saturated by top-5 retrieval; on TechQA, lexical / dense baselines are within a few percent of each other and Corpus2Skill pays a $30\times$ cost premium for no quality gain. The pattern is consistent: when the bottleneck for a corpus is retrieval recall rather than topic disambiguation, hierarchical navigation adds turns and tokens without helping. We therefore position Corpus2Skill as a complementary primitive — strongly preferred for curated, single-domain enterprise knowledge bases (the deployment target this paper is motivated by), with flat retrieval or RAPTOR retained as the right default for open-domain or homogeneous pools.
Competitive cases.
CovidQA and FinQA sit between the wins and the failures. On CovidQA the lexical metrics agree to within 0.02 F1 across all five methods. On FinQA, Corpus2Skill matches the best baseline on lexical n-gram quality (BLEU 0.132 and R-2 0.217 are both highest, F1 within 0.02 of Dense), confirming that financial reports still admit a useful by-firm and by-statement-section hierarchy; the cost premium ($\sim 80\times$ over Dense) rather than the quality is what makes it a poor fit. We treat FinQA and CovidQA as ties rather than losses in our headline count.
Failure modes exposed by the generalization sweep.
The CUAD and TatQA results are the two cleanest cases in our evaluation where the design hypothesis behind Corpus2Skill — that an LLM-written hierarchy provides useful navigation signal — does not hold. CUAD answers are long verbatim quotes from a single contract clause, which token-F1 penalizes for any paraphrase; TatQA is structurally so flat that the cluster summarizer has nothing discriminative to surface. Both motivate concrete follow-on work — span-level extraction tools at the leaf level for the former, table-aware schema-based clustering for the latter — discussed alongside other limitations in Section 5.
4.5 Ablation Studies
We conduct three ablation studies on the WixQA anchor to characterize how compile-time and serve-time design choices affect the quality-cost tradeoff.
Cluster structure.
Table 4: Ablation: cluster structure (200 queries, prompt caching on). All three shapes land within $\approx$2 F1 / 2 Fact points; the default shape is cheapest.
[…]
Narrow (p=5p{=}5, 3 skills, compact) 0.457 0.734 0.649 0.101
Default (p=10p{=}10, 6 skills) 0.468 0.739 0.673 0.089
Wide (p=20p{=}20, 2 skills) 0.469 0.749 0.666 0.096
We vary the branching ratio $p$ and top-cluster count $K$ to produce hierarchies of different shapes (Table 4). With p=5p{=}5, the narrow branching ratio creates a 4-level hierarchy with 3 top-level skills; with p=20p{=}20 we obtain a wide-shallow tree with only 2 very broad top-level skills that each span roughly 3,000 documents. With deeper hierarchies or larger corpora, the file count per skill can exceed the API’s 200-file limit. We introduce a compact mode that merges near-leaf INDEX.md files into their parent: instead of creating a separate subdirectory for each lowest-level group, the parent’s INDEX.md contains labeled sections listing each sub-group’s documents inline. This reduces file count by up to 80% while preserving all navigational information. All three shapes land within $\approx$2 F1 points and $\approx$2 Factuality points of each other (F1 0.457 / 0.468 / 0.469, Factuality 0.734 / 0.739 / 0.749); the default shape is cheapest at \$0.089 per query, with narrow 14% more expensive (\$0.101) and wide 8% more expensive (\$0.096). The takeaway is that once the per-level LLM-written summaries are faithful, the agent is robust to the tree’s branching factor: wider trees with fewer, broader top-level skills still navigate effectively because INDEX.md files at deeper levels restore the needed specificity, and narrower trees reduce the chance of top-level routing errors but add one extra navigation step per query. These results corroborate our main finding that the quality of the summaries at each level, not the precise shape of the tree, is the primary driver of navigation quality.
Agent exploration budget.
Table 5: Ablation: agent exploration budget (200 queries, prompt caching on). Quality is stable across budgets, indicating effective hierarchy organization.
[…]
5 0.458 0.741 0.657 0.087
10 0.464 0.737 0.670 0.095
20 0.468 0.739 0.673 0.089
We vary the maximum allowed interaction rounds while keeping the default skill tree (Table 5). The exploration budget has minimal impact on quality: even a budget of 5 achieves F1 of 0.458, just 2% below the full-budget configuration, and Factuality is nearly identical across variants (0.737–0.741). The hierarchical structure is sufficiently well-organized that the agent rarely needs extended exploration. Budget-5 also happens to be the cheapest setting (\$0.087 vs. \$0.089 for budget-20), because shorter traces generate and cache fewer intermediate turns; this makes max_turns=5 an attractive default for cost-sensitive deployments.
Serving LLM choice.
Table 6: Ablation: serving LLM choice (200 queries, prompt caching on). Haiku retains 91% of Sonnet’s F1 and matches its Context Recall at 71% of the cost.
[…]
Claude Sonnet 4.6 0.468 0.739 0.673 0.089
Claude Haiku 4.5 0.424 0.662 0.698 0.063
We swap only the serving model while keeping the same compiled skill tree (Table 6). Swapping Sonnet 4.6 for the smaller and cheaper Haiku 4.5 reduces per-query cost from \$0.089 to \$0.063 (a 29% reduction) while retaining 91% of Sonnet’s F1 (0.424 vs. 0.468) and 90% of its Factuality (0.662 vs. 0.739). Context Recall actually increases from 0.673 to 0.698, suggesting that Haiku navigates the hierarchy more thoroughly—retrieving more relevant documents—even if its answer synthesis is slightly less polished. At \$0.063 per query, Haiku offers a compelling cost-quality tradeoff for budget-constrained deployments. This also demonstrates that the compiled skill tree is robust to the choice of serving model: the quality of the hierarchy, not the sophistication of the navigator, is the primary driver of retrieval performance.
5 Limitations
Cost.
The per-query cost of Corpus2Skill is dominated by input tokens: the Skills API includes navigation file content in every call, so even relatively small SKILL.md and INDEX.md files accumulate significant token counts across turns. Enabling Anthropic ephemeral prompt caching on the navigation context reduces \$/query from \$0.172 to \$0.089—close to the Agentic baseline—since roughly 70% of the prompt content is reused across consecutive turns and queries; we use this cached configuration throughout our main results. Corpus2Skill is nonetheless best suited for high-value queries where answer quality justifies the cost premium, such as complex multi-hop support questions or compliance-critical lookups. For high-volume, low-stakes queries, single-shot retrieval or RAPTOR remain cost-effective.
API constraints.
The framework relies on Anthropic’s Skills API, which imposes limits of 8 skills per request, 200 files per skill, and 30 MB per skill. These constraints shape the hierarchy design: the 8-skill limit caps top-level clusters, and the 200-file budget constrains tree depth before requiring compact merging (Section 4.5). The WixQA corpus fits comfortably (6 skills, $\leq$133 files each), but larger corpora would require compact merging or multi-level partitioning. As Skills APIs evolve and constraints relax, even narrower hierarchies could improve navigation precision.
Hard single-path clustering.
Each document is assigned to exactly one cluster path, a deliberate choice that enables filesystem materialization and deterministic navigation. However, documents that span multiple topics (e.g., billing and subscriptions) can only appear under one branch, creating a routing bottleneck when queries target the other topic. Our failure analysis (Appendix G) confirms that top-level routing errors account for the majority of failures. Soft or multi-parent assignment could mitigate this at the cost of duplicated content and more complex navigation; we leave this exploration to future work.
Compilation.
The one-time compilation costs \$5–\$10 in LLM calls plus 6.5 minutes of wall-clock time, negligible for enterprise deployments handling thousands of daily queries. The pipeline is fully automated and deterministic given fixed random seeds, but does not yet support incremental updates—adding documents requires recompilation.
Corpora where corpus navigation does not help.
Our RAGBench generalization study (Section 4.4) makes the failure modes explicit. Corpus2Skill loses to flat retrieval on three structurally distinct corpus shapes. (1) Long extractive documents: CUAD legal contracts ($\approx$7,500 characters per doc, with answers being verbatim quotes from a single clause) and TechQA technotes (3,740 characters covering multiple symptom categories) violate the assumption that one document fits cleanly under one branch. (2) Homogeneous tabular corpora: TatQA is essentially identical SEC-style financial-statement tables, so the level-1 cluster summaries collapse onto near-duplicate labels and the navigator cannot disambiguate top-level branches; this case yields the worst quality and cost in our evaluation. (3) Open-domain pools where retrieval recall is the bottleneck: on CovidQA short research abstracts and ExpertQA cross-disciplinary scientific QA, hierarchical navigation adds turns and tokens without improving over what dense retrieval already surfaces in a single shot. We see these as honest scope limits of the design hypothesis behind Corpus2Skill rather than implementation issues. Three natural follow-on directions are clear from this analysis: a span-level retrieval head attached at the leaf level so the agent can quote verbatim from long documents (addressing CUAD), table-aware schema clustering that groups financial-statement rows by line-item / period / segment rather than embedding similarity (addressing TatQA), and hybrid routing that lets a lightweight classifier decide per-query whether to enter the skill tree or fall back to dense retrieval. We are exploring this last design under the working name CorpusForge in subsequent work.
6 Related Work
Cluster-based document organization.
Scatter/Gather (Cutting et al., 1992) introduced the paradigm of navigating a pre-computed cluster organization rather than matching by similarity, grounded in the cluster hypothesis (van Rijsbergen, 1979; Liu and Croft, 2004). Where Scatter/Gather required users to manually select and refine clusters in an interactive loop, Corpus2Skill delegates this navigation entirely to an LLM agent that reasons about cluster summaries at each level.
Hierarchical summarization for retrieval.
RAPTOR (Sarthi et al., 2024) recursively clusters and summarizes text chunks into a tree, retrieving via layer-by-layer traversal or collapsed-tree vector search. Corpus2Skill shares the compile-time embed-cluster-summarize loop but materializes the hierarchy as navigable files and replaces fixed traversal with agentic navigation. GraphRAG (Edge et al., 2024) builds a knowledge graph with community-level summaries; unlike Corpus2Skill, it requires a graph database at serve time. HiRAG (Huang et al., 2025) augments RAG with a hierarchical knowledge graph that interleaves entity and community levels; it still relies on graph retrieval at query time rather than agent-driven file navigation.
Retrieval-augmented generation.
RAG (Lewis et al., 2020) grounds LLM responses in retrieved passages, with dense passage retrieval (Karpukhin et al., 2020) and late-interaction models (Santhanam et al., 2022) pushing retrieval quality. Self-RAG (Asai et al., 2024) teaches the model to critique its own retrieval; IRCoT (Trivedi et al., 2023) interleaves retrieval with chain-of-thought reasoning for multi-hop questions, using intermediate reasoning steps to formulate better queries. Recent surveys (Gao et al., 2024; Singh et al., 2025) organize RAG along axes of granularity, timing, and agentic control. Corpus2Skill occupies a distinct point on this spectrum: rather than improving retrieval, it replaces embedding-based search with agent-driven file navigation over a pre-compiled hierarchy.
Structure-aware and agentic RAG.
StructRAG (Li et al., 2025) dynamically restructures retrieved information into tables, graphs, or catalogues at inference time, letting the LLM reason over structured views. BookRAG (Wang et al., 2025) builds a hierarchical index mirroring a document’s table of contents and uses an agent-based planner to query it. A-RAG (Du et al., 2026) exposes hierarchical retrieval interfaces (keyword, semantic, chunk-read) to an LLM agent, and SPD-RAG (Akay et al., 2026) assigns a dedicated sub-agent per document in a multi-agent pipeline. NaviRAG (Dai et al., 2026) builds a hierarchical view of documents and uses an LLM agent to actively navigate it, detecting information gaps and retrieving at the appropriate granularity. HCAG (Wu and Deng, 2026) constructs a multi-resolution knowledge base offline and performs hierarchical retrieval with a multi-agent discussion at query time. All of these systems retain some form of retrieval or search infrastructure at serve time. Corpus2Skill differs by materializing the hierarchy as a static filesystem of skill files and navigating it with a single agent using file browsing and document lookup rather than embedding-based search.
Tool-use agents.
Toolformer (Schick et al., 2023) and ReAct (Yao et al., 2023) established that LLMs can autonomously select and use external tools. Corpus2Skill extends this paradigm with domain-specific, read-only tools (file browsing and document lookup) over a static hierarchy, making navigation deterministic and reproducible given the same compiled skill tree.
Agent skill construction.
Voyager (Wang et al., 2023) builds a skill library from successful action sequences as code; SkillX (Wang et al., 2026) and EvoSkill (Alzubi et al., 2026) distill agent trajectories into reusable knowledge. These create skills from trajectories, whereas Corpus2Skill creates skills from document corpora via clustering and summarization, a fundamentally different input modality.
Positioning.
Corpus2Skill combines offline hierarchical clustering, materialization as navigable skill files, and LLM-driven top-down navigation into a system that replaces embedding-based retrieval with agent-driven file browsing at serve time. While NaviRAG and HCAG share the motivation of active hierarchical navigation, Corpus2Skill is distinguished by its filesystem-based skill representation, progressive disclosure via SKILL.md/INDEX.md files, and reliance on a standard code-execution tool rather than custom retrieval APIs. Table 8 in the Appendix provides a detailed feature comparison.
7 Conclusion
Corpus2Skill demonstrates that giving an LLM agent a structured, browsable overview of a document corpus leads to higher-quality answers than providing it with retrieved passages alone. The key insight is that modern LLM agents are capable navigators when given the right map: a pre-compiled hierarchy of summaries and topic clusters enables informed, top-down exploration that exploits the agent’s reasoning abilities rather than treating it as a passive consumer of search results. Because the agent reads summaries describing each branch, it can make informed routing decisions and combine evidence across branches when queries span multiple topics—an advantage that flat retrieval, which offers no corpus overview, cannot match. With Anthropic ephemeral prompt caching enabled on the navigation context, Corpus2Skill matches the Agentic baseline on per-query cost (\$0.089 vs. \$0.088) while still delivering the highest quality on every metric, and it suggests a complementary design point for knowledge-grounded systems: invest computation at compile time to make the corpus navigable, rather than at query time to make it searchable. Hierarchy depth grows logarithmically ($O(\log_{p}N)$), so the framework scales to larger corpora.
Future work includes evaluation on more domains, incremental compilation, and multi-parent soft clustering to handle queries that span multiple topics.
Impact Statement
Corpus2Skill is designed for enterprise knowledge-grounded QA, a setting where incorrect answers can lead to real-world harm (e.g., wrong compliance guidance or billing instructions). Like all LLM-based systems, it can produce confidently wrong answers when the agent navigates to irrelevant documents or misinterprets retrieved content. Hierarchical summaries produced during compilation may also expose corpus structure that organizations consider sensitive. Hard clustering means that documents spanning multiple topics are routed to a single branch, creating blind spots that may disproportionately affect certain query types. Finally, because compilation is a batch process, the system cannot reflect real-time corpus updates, risking stale or outdated answers until recompilation occurs. We recommend human-in-the-loop verification for high-stakes deployments.
Appendix A Compilation Details
Top-level skill labels.
The 6 skills produced by the default configuration (p=10p{=}10, K=7K{=}7) on the WixQA corpus are shown in Table 7.
Table 7: Top-level skills produced by Corpus2Skill on WixQA.
Skill Label Docs Coverage
business-tools 931 Scheduling, CRM
commerce 1,513 Store, payments
content-media 1,074 Blog, galleries
platform-guide 1,161 Editor, domains
marketing-seo 553 SEO, analytics
platform-overview 989 Members, apps
Embedding.
Documents are embedded using Qwen/Qwen3-Embedding-0.6B, a 622M-parameter model producing 1024-dimensional embeddings. We use the document prompt template with batch size 32. This model is chosen for its strong performance-to-size ratio and ability to run on CPU.
Summarization.
The compilation pipeline uses direct LLM API calls (not an agent loop) for summarization and labeling. Level-1 summaries condense up to 15 document texts (600 characters each); deeper-level summaries condense up to 20 child summaries (300 characters each). Summarization is performed concurrently using async calls with a semaphore of 20 for batch efficiency.
File counts.
The compilation produces 665 navigation files (SKILL.md + INDEX.md) and a documents.json file storing the full text of all 6,221 documents (13 MB). Total output size is 16 MB.
Document-level summarization prompt.
Used at the lowest level to summarize a cluster of raw documents.
You are summarizing a cluster of related
documents from a knowledge base. Write a
2-3 sentence summary that captures:
1. The common TOPIC area these documents
   cover
2. The types of QUESTIONS these documents
   answer
3. Key TERMS or features mentioned across
   documents
Be specific and concrete. Name actual
features, products, or processes.
Documents ({N} total, showing up to 15):
--- Document 1 ---
{doc_text[:600]}
--- Document 2 ---
{doc_text[:600]}
Summary:
Cluster-level summarization prompt.
Used at higher levels to summarize a group of sub-cluster summaries into a broader overview.
You are summarizing a level-{L} grouping
of {N} sub-groups from a knowledge base.
Each sub-group already has a summary below.
Write a 2-3 sentence overview that captures:
1. The broad DOMAIN these sub-groups cover
2. The range of TOPICS within this domain
3. What types of user QUESTIONS this group
   can answer
Be specific -- name the main product areas,
features, or workflows.
Sub-group summaries:
- Sub-group 1: {summary[:300]}
- Sub-group 2: {summary[:300]}
Overview:
Labeling prompt.
Used to generate a filesystem-safe directory name from a cluster summary. The response is post-processed: lowercased, non-alphanumeric characters replaced with hyphens, truncated to 50 characters.
Generate a short (2-5 word) filesystem-safe
label for this cluster. Use lowercase,
hyphens instead of spaces. No quotes.
Summary: {summary[:500]}
Appendix B Skill Structure Examples
We show the full content of representative navigation files from the WixQA compilation (default configuration: p=10p{=}10, K=7K{=}7). These files are what the agent reads during navigation.
Root-level SKILL.md.
This file sits at the top of the skill-01-wix-commerce-monetization skill directory, the largest skill covering 1,513 documents across 16 subgroups. The agent reads this file in a single tool call and uses the subgroup summaries to decide which branch to explore.
name: skill-01-wix-commerce-monetization
description: >
  This group covers the full Wix commerce
  and monetization ecosystem, spanning every
  major product vertical and workflow
  available to Wix merchants and business
  owners. Topics range from storefront and
  product management (Wix Stores,
  Restaurants, POS, Print/Fulfillment) to
  financial operations (payments, taxes,
  refunds, invoices, pricing plans,
  subscriptions, and promotional tools like
  coupons and gift cards)...
level: 3
num_documents: 1513
## Overview
This group covers the full Wix commerce and
monetization ecosystem, spanning every major
product vertical and workflow. Topics range
from storefront and product management (Wix
Stores, Restaurants, POS, Print/Fulfillment)
to financial operations (payments, taxes,
refunds, invoices, pricing plans,
subscriptions, and promotional tools), as
well as loyalty programs, partner ecosystems,
and learning/course delivery via Wix Online
Programs. Users can get answers about setting
up and running an online or in-person
business on Wix.
## Contents
### Sub-groups (directories)
Read the INDEX.md in each sub-group to
understand what it covers.
- **group-00-wix-restaurants-platform/**
  (159 docs): Covers the Wix Restaurants
  platform in its entirety, spanning both
  the original and newer versions of Wix
  Restaurants Orders and Menus apps.
- **group-01-wix-payments-ecosystem/**
  (87 docs): Covers Wix’s payment ecosystem,
  spanning payment processing, account
  management, and financial operations.
- **group-02-wix-pos-systems/** (110 docs):
  Covers Wix’s Point of Sale (POS)
  ecosystem, both tablet-based Retail POS
  and smartphone-based Mobile POS.
- **group-03-wix-stores-product-management/**
  (99 docs): Covers Wix Stores product and
  storefront management, from setup to
  customization.
- **group-04-wix-stores-shipping-config/**
  (57 docs): Covers shipping, pricing, and
  checkout configuration.
- **group-05-wix-payment-providers/**
  (289 docs): Covers payment processing and
  provider integration, both first-party
  and third-party gateways.
- **group-06-wix-subscription-management/**
  (133 docs): Covers subscription and pricing
  plan management across Premium, Studio,
  Editor X, Ascend, and App Market plans.
- **group-07-wix-print-fulfillment/**
  (55 docs): Covers print and fulfillment-
  related commerce features.
- **group-08-wix-commerce-fulfillment/**
  (74 docs): Covers order management,
  fulfillment, shipping, and inventory.
- **group-09-wix-online-programs/**
  (63 docs): Covers Wix Online Programs for
  building and delivering online courses.
- **group-10-refund-cancellation-policies/**
  (25 docs): Covers refund and cancellation
  policies across Wix products.
- **group-11-wix-pricing-plans/**
  (113 docs): Covers Wix Pricing Plans and
  monetization features lifecycle.
- **group-12-wix-partner-loyalty-programs/**
  (75 docs): Covers partner, loyalty, and
  rewards ecosystems.
- **group-13-taxes-fees-config/**
  (53 docs): Covers tax configuration,
  tips, and fees/surcharges.
- **group-14-wix-billing-documents/**
  (87 docs): Covers financial document
  management: invoices, quotes, proposals,
  receipts.
- **group-15-wix-payments-promotions/**
  (34 docs): Covers promotional tools:
  discount coupons, gift cards, and saved
  payment methods.
Mid-level INDEX.md.
This file sits one level below the root, inside group-01-wix-payments-ecosystem/, covering 87 documents across 10 leaf groups. The agent reads this after deciding from the SKILL.md above that payments are relevant.
name: wix-payments-ecosystem
description: >
  Covers Wix’s payment ecosystem, spanning
  payment processing, account management,
  and financial operations across Wix
  Payments and related commerce tools.
level: 2
num_documents: 87
## Overview
These sub-groups collectively cover Wix’s
payment ecosystem, spanning payment
processing, account management, and
financial operations across Wix Payments
and related commerce tools (Wix Stores,
Bookings, Restaurants, and Forms). Topics
range from setting up and verifying Wix
Payments accounts, managing payouts and
transactions, handling fraud, chargebacks,
and restricted products, to workflows like
collecting donations, processing manual
card entries, and understanding split
payment limitations.
## Contents
### Sub-groups (directories)
Read the INDEX.md in each sub-group to
understand what it covers.
- **group-00/** (6 docs): Wix account
  billing, payments management, and
  role-based permissions.
- **group-01/** (14 docs): Wix Payments
  payment processing, payout management,
  and account administration.
- **group-02/** (15 docs): Setup, mgmt,
  and compliance requirements of Wix
  Payments accounts.
- **group-03/** (7 docs): Accepting and
  managing online donations through Wix.
- **group-04/** (8 docs): Payment
  flexibility limitations across Wix
  products (split, defer, partial pay).
- **group-05/** (5 docs): Fraud prevention
  and payment security within Wix Payments.
- **group-06/** (8 docs): Wix Payments
  dashboard and transaction management.
- **group-07/** (12 docs): Credit and debit
  card payment processing within Wix.
- **group-08/** (4 docs): Payment provider
  compliance and restricted/prohibited
  products.
- **group-09/** (8 docs): Chargebacks and
  fees within Wix Payments.
Leaf-level INDEX.md.
This file sits at the bottom of the hierarchy, inside group-05/ under the payments ecosystem. It lists individual document IDs with titles. The agent reads this to select which documents to retrieve via get_document.
name: L1-C298
description: >
  Fraud prevention and payment security
  within Wix Payments, addressing merchant-
  facing and customer-facing concerns
  related to credit card transactions.
level: 1
num_documents: 5
## Overview
These documents cover fraud prevention and
payment security within Wix Payments,
addressing both merchant-facing and
customer-facing concerns related to credit
card transactions. They answer questions
about how to identify and respond to
suspicious or declined payments, understand
liability for fraudulent chargebacks, and
interpret unusual bank statement activity.
Key terms: liability shift, chargeback
mechanisms, chip card readers, suspicious
payment flags, fraud detection models.
## Contents
### Documents (5 items)
Use get_document tool with the doc_id
to read full content.
- ‘56ae7642b2a8c2ea‘: Best Practices for
  Accepting Card Payments
- ‘448d7947c5a22489‘: Wix Payments:
  About the ’Liability Shift’
- ‘bb73d9cb90385f12‘: Wix Payments:
  Declined Payment Due to Suspected Fraud
- ‘c3b01aac19482333‘: Wix Payments:
  About Suspicious Payments
- ‘d22b8cf6e599e0b8‘: Double Charge
  Appearing on Your Bank Statement
Appendix C Compilation Prompts
The compilation pipeline uses three LLM prompts, each issued as a single-turn API call (no agent loop or tool use).
Document-level summarization.
Used at the lowest level to summarize a cluster of raw documents.
You are summarizing a cluster of related
documents from a knowledge base. Write a
2-3 sentence summary that captures:
1. The common TOPIC area these documents
   cover
2. The types of QUESTIONS these documents
   answer
3. Key TERMS or features mentioned across
   documents
Be specific and concrete. Name actual
features, products, or processes.
Documents ({N} total, showing up to 15):
--- Document 1 ---
{doc_text[:600]}
--- Document 2 ---
{doc_text[:600]}
Summary:
Cluster-level summarization.
Used at higher levels to summarize a group of sub-cluster summaries into a broader overview.
You are summarizing a level-{L} grouping
of {N} sub-groups from a knowledge base.
Each sub-group already has a summary below.
Write a 2-3 sentence overview that captures:
1. The broad DOMAIN these sub-groups cover
2. The range of TOPICS within this domain
3. What types of user QUESTIONS this group
   can answer
Be specific -- name the main product areas,
features, or workflows.
Sub-group summaries:
- Sub-group 1: {summary[:300]}
- Sub-group 2: {summary[:300]}
Overview:
Labeling.
Used to generate a filesystem-safe directory name from a cluster summary. The response is post-processed: lowercased, non-alphanumeric characters replaced with hyphens, truncated to 50 characters.
Generate a short (2-5 word) filesystem-safe
label for this cluster. Use lowercase,
hyphens instead of spaces. No quotes.
Summary: {summary[:500]}
Appendix D Serving System Prompt
The serving agent receives the following system prompt:
# Corpus-Grounded Support Agent
#   (Hierarchical Navigation)
You are a knowledge agent that answers
questions by navigating a hierarchical
skill directory. You explore a structured
file tree where documents are organized
into topic clusters.
## Hard Rules
- Every factual claim must trace to a
  document you retrieved via get_document.
- SKILL.md files are NAVIGATION AIDS --
  they tell you where to look, not what
  to say.
- Never fabricate steps, URLs, prices, or
  specifics not found in documents.
- Never guess. If you cannot find relevant
  content after thorough exploration,
  say so.
## Navigation Strategy
Your skills directory has this structure:
  skill-$X$-topic/
    SKILL.md        <- top-level summary
                       + index of contents
    group-$Y$-subtopic/
      INDEX.md      <- summary + index of
                       sub-contents
      group-ZZ/
        INDEX.md    <- summary + document
                       ID listings
Skill names and descriptions are already
available to you. Follow this workflow:
1. Read the SKILL.md of the 1-2 most
   relevant skills for your query.
2. Drill into the most relevant sub-group:
   Read its INDEX.md.
3. At the leaf level, INDEX.md lists
   document IDs with brief titles. Pick
   the most relevant document IDs.
4. Call get_document with each relevant
   doc_id to retrieve the full text.
5. Read at least one full document before
   answering.
## Tools
- Code execution: Use ‘ls‘ and ‘cat‘ to
  navigate the skills hierarchy.
- get_document(doc_id): Retrieve the full
  text of a document by its ID. The doc_id
  values are listed in leaf-level INDEX.md
  files.
## Answer Format
- First sentence = direct answer.
  No preamble.
- Factual questions: 1-3 sentences
  (~80 words max).
- Procedural questions: numbered steps
  only (~150 words max).
- Plain text. No bold, headers, or
  dividers.
- One approach only. Do not present
  alternatives.
- Never add "contact support" or closing
  remarks.
Appendix E Full Navigation Traces
We present complete navigation traces with raw tool calls for four queries from the evaluation set.
Trace 1: Direct descent (Factuality 1.0, CtxR 1.0).
Query: “I need to switch my business type from sole prop to LLC in order to use an EIN.”
[Context: agent sees 6 skill descriptions]
--- Turn 1 ---
Agent thought: The query is about changing
business type for Wix Payments. The skill
"wix-commerce-monetization" covers payments,
taxes, and business operations.
Tool: text_editor_code_execution
  view skill-01-wix-commerce-monetization/
    SKILL.md
Result: 10 subgroups listed, including:
  - group-01-wix-payments-ecosystem/
    (87 docs): "payment processing,
    account management, financial
    operations"
--- Turn 2 ---
Tool: text_editor_code_execution
  view group-01-wix-payments-ecosystem/
    INDEX.md
Result: 8 leaf groups. Agent identifies
  group-05 covering Wix Payments account
  setup and verification.
Tool: text_editor_code_execution
  view group-05/INDEX.md
Result: 12 documents listed, including:
  - d56cc797742e: "Changing Your Wix
    Payments Account Type"
--- Turn 3 ---
Tool: get_document
  {"doc_id": "d56cc797742e7d0f"}
Result: Full article text (1,247 chars)
  describing that account type cannot be
  changed directly; must contact Wix
  Customer Care.
--- Answer ---
"You cannot change your account type
directly within Wix Payments. To switch
from sole proprietorship to LLC, contact
Wix Customer Care with your EIN and LLC
formation documents ready."
Navigation: 6221 -> 1513 -> 87 -> 12 -> 1
Trace 2: Multi-document retrieval (Factuality 0.8, CtxR 1.0).
Query: “How do I sync the hotel app with my calendars iCal link to allow visitors to book available dates?”
[Context: agent sees 6 skill descriptions]
--- Turn 1 ---
Agent thought: Hotels and scheduling fall
under "wix-business-tools" which covers
scheduling, CRM, and business management.
Tool: text_editor_code_execution
  view skill-00-wix-business-tools/SKILL.md
Result: 8 subgroups listed, including:
  - group-02-wix-hotels-platform/
    (82 docs): "hotel management, bookings,
    reservations, calendar sync"
--- Turn 2 ---
Tool: text_editor_code_execution
  view group-02-wix-hotels-platform/INDEX.md
Result: 7 leaf groups. Agent identifies
  group-05 covering external calendar
  integration and iCal sync.
Tool: text_editor_code_execution
  view group-05/INDEX.md
Result: 17 documents listed, including:
  - 5f02fc2f8937: "Importing Reservations
    from External Calendar (iCal)"
  - 741830b917b4: "Connecting Airbnb iCal
    to Your Site"
--- Turn 3 ---
Tool: get_document
  {"doc_id": "5f02fc2f8937d16a"}
Result: Step-by-step iCal import guide
  for original Wix Hotels.
Tool: get_document
  {"doc_id": "741830b917b47d07"}
Result: Airbnb iCal connection guide for
  Wix Hotels by HotelRunner.
--- Answer ---
"For original Wix Hotels: Go to Hotels
tab > Reservations > iCal, select room
type, paste iCal URL, click Save. For
HotelRunner: go to dashboard, navigate
to channel connections, add Airbnb iCal
link for each property..."
Navigation: 6221 -> 931 -> 82 -> 17 -> 2
Trace 3: Mobile quick action (Factuality 1.0, CtxR 1.0).
Query: “I need to temporarily remove the quick action button from the mobile view.”
--- Turn 1 ---
Agent thought: Mobile site editing falls
under "wix-platform-guide" which covers
Editor, site building, and customization.
Tool: text_editor_code_execution
  view skill-03-wix-platform-guide/SKILL.md
Result: 10 subgroups. Agent picks
  "wix-mobile-site-building" covering
  mobile editor features.
--- Turn 2 ---
Tool: text_editor_code_execution
  view wix-mobile-site-building/INDEX.md
Result: Leaf groups covering mobile UI
  elements. Agent finds group with docs
  on Quick Action Bar.
Tool: text_editor_code_execution
  view group-03/INDEX.md
Result: 8 documents including:
  - bce5e3d829fe: "Adding and Setting Up
    Your Quick Action Bar"
--- Turn 3 ---
Tool: get_document
  {"doc_id": "bce5e3d829fe6eeb"}
Tool: get_document
  {"doc_id": "186ad4a2da9f1f30"}
--- Answer ---
"There is no hide option, only deletion.
Right-click the Quick Action Bar in the
mobile Editor and select Delete. To re-add
later, go to Mobile Tools and click Add."
Navigation: 6221 -> 1161 -> 142 -> 8 -> 2
Trace 4: Cross-branch exploration (Factuality 1.0, CtxR 0.6).
Query: “I want to change the currency for my course.”
[Context: agent sees 6 skill descriptions]
--- Turn 1 ---
Agent thought: "Course" suggests Online
Programs, and "currency" is a payments/
billing topic. The skill "wix-commerce-
monetization" covers both online programs
and payment configuration.
Tool: text_editor_code_execution
  view skill-01-wix-commerce-monetization/
    SKILL.md
Result: 15 subgroups listed, including:
  - group-09-wix-online-programs/
    (107 docs): "online courses, program
    management, pricing, delivery"
  - group-14-wix-billing-documents/
    (68 docs): "currency settings, invoice
    automation, financial documents"
--- Turn 2 ---
Agent thought: Start with Online Programs
since the query is about course currency.
Tool: text_editor_code_execution
  view group-09-wix-online-programs/INDEX.md
Result: 5 leaf groups. Picks group-03
  covering program settings (18 docs).
Tool: text_editor_code_execution
  view group-03/INDEX.md
Result: 18 documents including:
  - 7a9dc6f2de90: "Online Programs:
    Editing Your Program Settings"
--- Turn 3 ---
Tool: get_document
  {"doc_id": "7a9dc6f2de90db61"}
Result: Program settings article. Mentions
  that currency is configured at the site
  level via Language & Region settings,
  not per-program.
Agent thought: Need the actual currency
configuration steps. Navigate to billing.
Tool: text_editor_code_execution
  view group-14-wix-billing-documents/
    INDEX.md
Result: 5 leaf groups. Picks group-00
  covering currency management (18 docs).
Tool: text_editor_code_execution
  view group-00/INDEX.md
Result: 18 documents including:
  - 29139bd7c379: "Setting Your Currency
    for Accepting Payments"
Tool: get_document
  {"doc_id": "29139bd7c379ff77"}
Result: Step-by-step currency configuration
  guide via Settings > Language & Region.
--- Answer ---
"To change the currency for your course,
update the site-level currency setting:
1. Go to Settings in your dashboard.
2. Click Language & Region.
3. Scroll to Currency and select your
   desired currency.
4. Click Save.
This applies to all Wix products including
Online Programs. Your currency must match
your payment provider’s currency."
Navigation paths:
  6221 -> 1513 -> 107 -> 18 -> 1 (programs)
  6221 -> 1513 ->  68 -> 18 -> 1 (billing)
This trace demonstrates cross-branch navigation: the agent first explored the online-programs subgroup to understand the relationship between courses and currency, then navigated to the billing-documents subgroup to find the authoritative currency configuration steps. The two branches within the same skill provided complementary information that neither alone would have answered the query fully.
Appendix F Positioning Comparison
Table 8: Positioning of Corpus2Skill relative to prior work.
Aspect Scatter/Gather RAPTOR GraphRAG Voyager/SkillX Corpus2Skill
Input Documents Text chunks Documents Trajectories Documents
Organization Flat clusters Hier. tree Knowledge graph Skill library Hier. skill tree
Navigation Human user Algorithm Graph search Vector sim. LLM agent
Serve infra Cluster index Vector index Graph DB Vector DB LLM only
Table 8 highlights three axes on which Corpus2Skill differs from prior work.
Input modality.
Scatter/Gather, RAPTOR, GraphRAG, and Corpus2Skill all start from documents or text chunks, but Voyager and SkillX build skills from agent trajectories—successful action sequences or distilled experience. Corpus2Skill is the only system that converts a static document corpus into the skill format typically reserved for procedural knowledge.
Navigation mechanism.
Scatter/Gather relies on human users to iteratively select clusters; RAPTOR and GraphRAG delegate navigation to fixed algorithms (vector similarity or graph traversal). Corpus2Skill is unique in delegating navigation to an LLM agent that can reason about which branch to explore, backtrack when a path is unproductive, and combine evidence across branches. This agent-driven navigation is what enables the trial-and-error behavior illustrated in Figures 4 and 5.
Serve-time infrastructure.
All prior systems require dedicated infrastructure at query time: a cluster index, a vector database, or a graph database. Corpus2Skill eliminates this dependency entirely—the only runtime component is the LLM itself, which receives pre-compiled skill files via the API. This “LLM only” property simplifies deployment but shifts the cost burden to input tokens, as discussed in Section 5.
Appendix G Failure Analysis
[…]
Navigation miss (38 queries).
The agent retrieves documents but none match the gold article, indicating it descended into the wrong branch of the hierarchy. For example, a query about reordering CMS collection items is routed to a general site-editing skill instead of the CMS-specific one; a query about blocking time slots in a staff schedule navigates to a generic booking group rather than the staff-management leaf. These misroutes occur at the top level of the hierarchy, where broad skill summaries fail to distinguish closely related topics.
Partial navigation (19 queries).
The agent’s retrieved set includes the gold document, yet scores remain low. In these cases the gold article is retrieved alongside several non-relevant documents, diluting the context provided to the answer synthesis step. The agent correctly identified the general topic area but over-retrieved from neighboring leaf groups.
Synthesis error (3 queries).
[…]
No documents retrieved (2 queries).
Two queries (“How do I clear my browser cache?” and a Gmail password reset question) resulted in zero retrieved documents. Both have gold articles in the corpus (Wix publishes help articles for these topics), but the agent could not locate a matching skill branch and correctly declined to answer rather than hallucinating.
Takeaway.
The dominant bottleneck is initial topic routing: 38 of 62 failures stem from the agent choosing the wrong top-level branch. The compact-mode ablation (Section 4.5), which produces narrower top-level clusters, partially mitigates this by providing finer-grained topic separation at the first navigation step.
Appendix H Baseline Implementation Details
All baselines use the same LLM (Claude Sonnet) for answer generation and the same evaluation prompts for fair comparison.
BM25.
We implement BM25 sparse keyword retrieval using the rank_bm25 library (Robertson and Zaragoza, 2009). Documents are tokenized and indexed at the full-article level. At query time, we retrieve the top-5 articles by BM25 score and pass them as context to a single LLM call.
Dense.
We encode all documents using Qwen3-Embedding-0.6B (Zhang et al., 2025), producing 1024-dimensional embeddings stored in a FAISS flat index (Johnson et al., 2019). We use the document prompt template for encoding articles and the query prompt template for encoding questions. At query time, we retrieve the top-5 nearest neighbors by cosine similarity and pass them as context to a single LLM call.
Hybrid.
We combine BM25 and Dense scores using Reciprocal Rank Fusion (Cormack et al., 2009) with k=60k{=}60 and equal weighting ($\alpha{=}0.5$). The fused ranking selects the top-5 passages, which are passed to a single LLM call.
RAPTOR.
We faithfully implement the RAPTOR algorithm (Sarthi et al., 2024): UMAP (McInnes et al., 2018) for dimensionality reduction, Gaussian Mixture Model (GMM) soft clustering (McLachlan and Peel, 2000) with BIC-based cluster count selection (Schwarz, 1978), and recursive summarization of clusters. We use the same embedding model (Qwen3-Embedding-0.6B) as the Dense baseline for consistency. At query time, collapsed tree retrieval searches across all hierarchy levels simultaneously, returning the top-5 passages from any level.
Agentic.
An LLM agent is given access to three search tools (search_bm25, search_dense, search_hybrid), each returning top-$k$ results. The system prompt instructs the agent to search first, then answer based on accumulated evidence, choosing BM25 for exact terms and product names, dense for conceptual queries, and hybrid as a default strategy. The agent can issue up to 10 rounds of iterative retrieval, reformulating sub-queries based on intermediate results. Evidence is accumulated across all turns; when the agent stops, it produces a structured JSON response containing the answer and curated top-5 references selected from the full evidence pool. For evaluation fairness, context recall is computed over the last 5 retrieved documents (matching the context window of single-shot baselines). This baseline tests whether agentic multi-turn retrieval over traditional search infrastructure can match navigation-based exploration.
Appendix I Evaluation Metrics
Lexical metrics.
Token F1 tokenizes both the predicted and gold answers into word tokens and computes the harmonic mean of token-level precision and recall. BLEU (Papineni et al., 2002) measures modified n-gram precision (up to 4-grams) with a brevity penalty to penalize overly short outputs. ROUGE-1 and ROUGE-2 (Lin, 2004) measure unigram and bigram recall respectively, capturing how much of the gold answer’s content appears in the prediction.
LLM-judged metrics.
Both Factuality and Context Recall use an LLM judge (Zheng et al., 2023) following the RAGAS framework (Es et al., 2024), scoring on a 1–5 scale normalized to 0–1. Factuality evaluates whether the claims in the generated answer are factually consistent with the gold answer, independent of the retrieved context. Context Recall evaluates whether the retrieved context contains sufficient information to answer the question, by checking how many essential claims in the gold answer can be attributed to the provided context.
Context Recall standardization.
A critical design decision is the standardization of Context Recall measurement across different retrieval paradigms (single-shot, multi-turn, and navigation-based). For fair comparison, all methods follow the same protocol: (1) collect the last 5 documents or passages retrieved by the method; (2) join them with separator tokens; (3) truncate to a uniform 8,000-character window. The LLM judge receives this standardized context. This ensures that single-shot baselines contribute their top-5 retrieved passages, Agentic contributes the last 5 documents from its multi-turn retrieval, and Corpus2Skill contributes the last 5 documents retrieved via get_document tool calls.
Cost metrics.
Input token count is the total number of tokens sent to the LLM across all turns. Per-query cost is computed using the model’s published per-token pricing (input and output tokens combined). For single-shot methods, cost reflects one LLM call; for multi-turn methods (Agentic, Corpus2Skill), it reflects all turns including tool-call reasoning.